OpenStudy (anonymous):

What is an equation that cannot be factored but can be solved using the quadratic formula. Example: x^2 - 2 + 1

OpenStudy (anonymous):

it's x^2 -1, so then =(x-1)(x+1) so , x=1,or x =-1////////pay attention to what you did write , but if you forgot to multiply -2 by x i mean if x^2 - 2x + 1 is your correct equation , then (x-1)^2=0 means that x=1

OpenStudy (anonymous):

I just need an equation that cannot be factored :/

OpenStudy (anonymous):

@SnuggieLad @ganeshie8